OpenStudy (hellomynameiskam):

Kathleen is a sales associate in a jewelry store. She earns $560 per week plus an 8% commission on sales. How much does she need to sell in a week to earn at least $700 that week?

OpenStudy (whpalmer4):

Kathleen's earnings are \(560+8\%*\text{sales}\) Set that equal to \(700\) and solve for the amount of sales.

OpenStudy (hellomynameiskam):

im confused @whpalmer4

OpenStudy (mathstudent55):

Kathleen earns $560 per week plus an amount that depends on her sales. Let's say Kathleen sells x amount of dollars in a week. The commission she will earn on the sales is 8% of x. Since 8% = 0.08, then 8% of x is 0.08x That means that for that week, she will earn 560 + 0.08x. Ok so far?

OpenStudy (hellomynameiskam):

yes.

OpenStudy (hellomynameiskam):

560 + 0.08x = 560.08 thats what i got

OpenStudy (mathstudent55):

No. She wants to earn 700, so we need to use 700. Also, she wants to earn at least $700. That means this problem needs an inequality.

OpenStudy (mathstudent55):

Since she wants to earn at least $700, that means the amount she earns must be greater than or equal to 700. She will earn 560 + 0.08x We want that amount to be greater than or equal to 700, so we setup an inequality: \(560 + 0.08x \ge 700\)

OpenStudy (mathstudent55):

The inequality above uses the unknown amount in sales, x, that she needs to sell to earn at least $700 in that week. Now you need to solve the inequality for x to find how much she needs to sell in that week.

OpenStudy (whpalmer4):

It's a fine point, but the amount she NEEDS to sell to earn at least $700 is given by \[700 = 560 + 0.08x\]Any sales beyond that value of \(x\) simply pad her income. There is only one value of \(x\) which is the minimum amount of sales she can make to earn $700 for the week. "How much does she need to sell" -> "What is the minimum value of sales"